OpenStudy (anonymous):

1+ sec^2x sin^2x = sec^2x Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[\begin{align*}1+\sec^2x\sin^2x&=1+\frac{\sin^2x}{\cos^2x}\\ &=1+\tan^2x\\ &=\sec^2x\end{align*}\] The last step is the Pythagroean identity, \[\sin^2x+\cos^2x=1~~\iff~~\tan^2x+1=\sec^2x~~\iff~~1+\cot^2x=\csc^2x\]

OpenStudy (anonymous):

ok, thank you :)